Shift summary DAY 12/13/99

Richard Ford (rickford@fnal.gov)
Mon, 13 Dec 1999 15:42:41 -0600 (CST)

Shift Summary for day shift, 12/13/99

On Shift: E. Turner, R. Ford

Hours of beam available at all intensities: 7.8
Hours of beam available at nominal intensity: 7.8
Hours of E799 data (special runs not included): 7.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

TeV PS water leak, NS1 LCW trip, spurious aborts

========================================================================
Runs ended this shift: 15211

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:
Beam intensity was mostly optimal. Some beam quality problems manifested
by out of limit scaler rates. Overall decent datataking until end of
shift with TeV PS water leak.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15211 07:30 Still getting SCALER/SEM errors; both high and
high and low limit. Called MCR to work on
spill quality.

07:56 HCC ineff. erros on channels 366,1767

08:00 Scaler errors much less frequent.

235-239 08:55 No Beam

09:00 Wire 84 of plane 9 has bad T0 offset

09:30 ET zeros Reg HV

10:25 PSH adjusted variable filter wheel for laser system
Laser light level increased by more than x2.

11:00 BobT and RickF look at rates and timing on new ACC4
counter. Results in log book.

11:55 Hogan wants be paged when beam is down 15min or more
so he can start a new run.

450 13:45 Scaler alarms go wild; ET notices epicure red light
Further investigation shows NM1U and NM1H tripped off.
Called MCR. They were unaware so we requested they
turn off beam until fixed.

459 13:52 Beam returns; LCW trip.

14:05 Spill structure gets bad. MCR informed.

515 15:07 Ch.13: Investigating TeV PS water leak.
Stop run 15211. Hogan calls within seconds of beam
disappearing as if he somehow knew. He is enroute.
Various special interests are organizing for accesses.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
CsI 2081 has a large gain drift(~5%), keep an eye on it.
HCC channels 2803 and 1767 were both highly inefficient
during the run, watch to make sure that they aren't getting worse,
experts have been notified.